### 15 Apr 2026 v2

# Hierarchical Reinforcement Learning with Augmented Step-Level Transitions for LLM Agents

Abstract
Large language model (LLM) agents have demonstrated strong capabilities in complex interactive decision-making tasks. However, existing LLM agents typically rely on increasingly long interaction histories, resulting in high computational cost and limited scalability. In this paper, we propose STEP-HRL, a hierarchical reinforcement learning (HRL) framework that enables step-level learning by conditioning only on single-step transitions rather than full interaction histories. STEP-HRL structures tasks hierarchically, using completed subtasks to represent global progress of overall task. By introducing a local progress module, it also iteratively and selectively summarizes interaction history within each subtask to produce a compact summary of local progress. Together, these components yield augmented step-level transitions for both high-level and low-level policies. Experimental results on ScienceWorld and ALFWorld benchmarks consistently demonstrate that STEP-HRL substantially outperforms baselines in terms of performance and generalization while reducing token usage. Our code is available at https://github.com/TonyStark042/STEP-HRL.
Hierarchical Reinforcement Learning with Augmented Step-Level Transitions for LLM Agents
Shuai Zhen1, Yanhua Yu1*, Roupei Guo2*, Nan Cheng2, Yang Deng3 1Beijing University of Posts and Telecommunications 2China Mobile Group Design Institute Co., Ltd 3Singapore Management University *Correspondence: yuyanhua@bupt.edu.cn, guoruopei@cmdi.chinamobile.com
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities as autonomous agents in sequential decision-making tasks, exhibiting sophisticated reasoning and planning abilities across diverse interactive environments (Wang et al., 2023; Yao et al., 2022; Li et al., 2022; Ren et al., 2025). To further enhance the effectiveness of autonomous agents, reinforcement learning (RL) offers a principled mechanism for enhancing agent decision-making capabilities  (Xu et al., 2024; Pang et al., 2024; Peiyuan et al., 2024). Unlike supervised approaches that rely solely on fixed demonstrations (Zeng et al., 2024a; Lin et al., 2023), RL enables agents to refine policy through environmental interaction and reward feedback, thereby discovering more effective strategies that generalize beyond training distributions.
Despite this progress, most LLM agents adopt a history-conditioned formulation, where policies are conditioned on increasingly long sequences of past observations and actions. This design choice is largely inherited from sequence-modeling perspectives: LLM agents are built on Transformer architectures (Vaswani et al., 2017), and recent RL formulations cast decision-making as trajectory or sequence prediction (Chen et al., 2021; Janner et al., 2021; Ni et al., 2023). While long histories can help infer latent states in partially observable environments, conflating long-horizon decision-making with long-context conditioning introduces fundamental limitations. Attention-based inference scales quadratically with context length, and unfiltered histories accumulate redundant or irrelevant information that can obscure decision-critical signals and degrade reasoning quality (Zhou et al., 2025; Cherepanov et al., 2023). Importantly, long-context conditioning is a modeling choice rather than a necessity of reinforcement learning.
Existing approaches primarily mitigate the symptoms of this formulation without revisiting its core assumption. Prior work compresses interaction histories (Zhou et al., 2025; Luo et al., 2024) or improves long-term credit assignment (Liu et al., 2025; Zhai et al., 2025; Xiong et al., 2024), but policies remain history-conditioned. Hierarchical reinforcement learning (HRL) introduces temporal abstraction and shows promise for LLM agents (Hu et al., 2025), yet current HRL methods still condition both high-level and low-level policies on accumulated interaction histories, inheriting the same long-context dependence they seek to alleviate.
To address the challenges discussed above, we propose STEP-HRL (Augmented Step-level Hierarchical Reinforcement Learning), which rethinks long-horizon LLM agents from a progress-based perspective. With completed high-level subtasks providing a global progress of overall task, STEP-HRL introduces an additional local progress module that accumulates subtask-relevant information at each timestep into a compact textual representation with controlled verbosity. The low-level policy conditions exclusively on the current subtask, observation and the distilled local progress, enabling step-level decision making with constant-sized inputs. Meanwhile, the local progress interacts with both the low-level and high-level policies, as well as with its own internal state, facilitating structured information transfer across hierarchical levels. We first perform behavior cloning on expert demonstrations to initialize policies, and then apply step-level offline RL for further optimization.
In summary, our contributions are as follows:
  • We propose STEP-HRL, a hierarchical framework that leverages a local progress module to enable policies to condition on single-step transitions for LLM agents, eliminating the need to condition on full interaction histories.
  • We propose a parameter-efficient two-stage training pipeline, where the high-level, low-level and local progress policies share a unified policy backbone, while being equipped with separate value networks for offline RL. The model is first initialized via behavior cloning and subsequently fine-tuned with step-level offline optimization.
  • Extensive experiments on the ScienceWorld and ALFWorld benchmarks demonstrate that our approach significantly improves both performance and generalization, validating the feasibility of step-level RL for LLM agents.
2 Problem Formulation
We formulate the agent operating in an interactive environment as a Partially Observable Markov Decision Process (POMDP), defined by the tuple $\langle\mathcal{C},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{T},\mathcal{R}\rangle$. Here, $\mathcal{C}$ denotes the instruction space specifying task goals, $\mathcal{S}$ is the latent environment state space, $\mathcal{A}$ is the action space, $\mathcal{O}$ is the observation space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ represents the state transition dynamics, and $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function. In the setting of LLM agents, $\mathcal{C},\mathcal{A},\mathcal{O}$ are expressed in natural language, while the environment state remains unobserved. Given an instruction $c\in\mathcal{C}$, the agent interacts with the environment with policy $\pi_{\theta}$, the policy parameters $\theta$ are initialized from a pretrained LLM. The objective is to optimize the policy to maximize the expected discounted return $J(\pi)=\mathbb{E}_{\pi}\!\left[\sum_{t}\gamma^{t}r_{t}\right].$
3 Method
Figure 1: (a): The pipeline of STEP-HRL. Local progress policy is responsible for producing a compact summary of local interaction history within each subtask. Specifically, the local progress policy $\pi^{p}_{\theta}$ depends on previous progress $p^{k}_{t-1}$, current subtask $g_{k}$, executed action $a^{k}_{t-1}$ and the resulting observation $o^{k}_{t}$ to the generate updated local progress $p^{k}_{t}$. The low-level policy $\pi^{l}_{\theta}$ combines $p^{k}_{t}$ with observation $o^{k}_{t}$ and subtask $g_{k}$ to generate primitive actions. When current subtask $g_{k}$ terminates, its final local progress $\hat{p}_{k}$ is forwarded to the high-level policy $\pi^{h}_{\theta}$. Conditioned on the task instruction $c_{i}$, completed subtasks $G_{k}$, final local progress $\hat{p}_{k}$ and the initial observation $o^{k+1}_{0}$ of next subtask, $\pi^{h}_{\theta}$ generates the subsequent subtask. (b): The structure of our model. Three different policies share the same parameters, but equipped with different critic network respectively for offline RL training.
3.1 Step-Level Transitions with Local Progress Modeling
Consider a commonly adopted history-conditioned RL formulation with hierarchical structures. Assume that all previous subtasks have been completed and a new subtask is to be generated. The global interaction history up to time $t$ is denoted as $\mathcal{H}_{t}=(c,o_{0},g_{0},a_{0},\ldots,a_{t-1},o_{t})$, which concatenates the task instruction with past observations, subtasks, and actions. Conditioned on $\mathcal{H}_{t}$, the high-level policy generates the next subtask:
$g_{k+1}\sim\pi^{h}_{\theta}(\cdot\mid\mathcal{H}_{t}).$ (1)
Given the $k$-th subtask $g_{k}$, the low-level policy operates at a finer temporal resolution. We denote the local interaction history as $h^{k}_{t}=(o^{k}_{0},a^{k}_{0},\ldots,o^{k}_{t})$, which records observations and actions accumulated during subtask $g_{k}$. Conditioned on the $g_{k}$ and $h^{k}_{t}$, the low-level policy produces primitive actions:
$a^{k}_{t}\sim\pi^{l}_{\theta}(\cdot\mid g_{k},h^{k}_{t}).$ (2)
The local history grows until the subtask terminates, after which the high-level policy is invoked again based on the updated global interaction history.
To enable step-level transitions, a key challenge is compactly representing both local and global interaction histories. Intuitively, the sequence of completed subtasks $G_{k}=(g_{0},g_{1},\ldots,g_{k})$ already serves as a concise summary of global task progress. Thus, the remaining problem is to compactly summarize the local interaction history within each subtask. To this end, we introduce a local progress policy $\pi_{\theta}^{p}$ to iteratively achieve this. At the beginning of the $g_{k}$, the local progress is initialized as $p^{k}_{0}=\varnothing$, reflecting the absence of subtask-local interaction history. The local progress is then updated at each subsequent timestep according to:
$p^{k}_{t}\sim\pi_{\theta}^{p}(\cdot\mid g_{k},a^{k}_{t-1},o^{k}_{t},p^{k}_{t-1}),\quad t>0$ (3)
This design encourages $\pi_{\theta}^{p}$ to selectively extract subtask-relevant information from the previous progress $p^{k}_{t-1}$ and integrate it with the last executed action $a^{k}_{t-1}$ and its resulting observation $o^{k}_{t}$, yielding an updated local progress $p^{k}_{t}$.
With the $p^{k}_{t}$ capturing subtask-relevant local interaction information, the low-level policy can make decisions based on the augmented step-level transition $(o^{k}_{t},p^{k}_{t},a^{k}_{t},\hat{r}^{k}_{t},o^{k}_{t+1},p^{k}_{t+1})$, where $\hat{r}^{k}_{t}$ is the intrinsic reward which equals $1$ if the current step successfully completes subtask $g_{k}$ and 0 otherwise. This formulation enables step-level decision making without relying on the full interaction history within each subtask:
$a^{k}_{t}\sim\pi^{l}_{\theta}(\cdot\mid g_{k},p^{k}_{t},o^{k}_{t}).$ (4)
It is worth noting that although $p^{k}_{t}$ already encodes information from current observation $o^{k}_{t}$, $o^{k}_{t}$ is typically most relevant for current action generation. To prevent $\pi^{p}_{\theta}$ from overlooking critical instantaneous information and to strengthen the sensitivity of $\pi^{l}_{\theta}$ to the current observation, we still explicitly include $o^{k}_{t}$ as an input to the low-level policy.
The local progress $p^{k}_{t}$ can also facilitate high-level subtask generation. If we restrict high-level policy $\pi^{h}_{\theta}$ to condition solely on the completed subtasks $G_{k}$, it does not observe the detailed low-level progress. In this setting, the local progress $p^{k}_{t}$ bridges this information gap. For simplicity, we denote the final local progress at the termination of subtask $g_{k}$ as $\hat{p}_{k}$. We pass the final progress $\hat{p}_{k-1}$ from the preceding subtask $g_{k-1}$ to the high-level policy. As a result, the step-level transition of high-level can be expressed as $(\hat{p}_{k-1},o^{k}_{0},g_{k},R_{k},\hat{p}_{k},o^{k+1}_{0})$, where $R_{k}=\sum_{t}r^{k}_{t}$ is the accumulated extrinsic environment reward during subtask $g_{k}$, and the high-level policy generates the next subtask according to:
$g_{k+1}\sim\pi^{h}_{\theta}(\cdot\mid c,G_{k},\hat{p}_{k},o^{k+1}_{0}).$ (5)
We adopt a parameter-efficient design across the three policies, $\pi^{h}_{\theta}$, $\pi^{l}_{\theta}$ and $\pi^{p}_{\theta}$ share the same parameters. This formulation facilitates efficient knowledge transfer across different decision levels, encourages consistent representations of task semantics and environment dynamics, and reduces the overall training and inference overhead. As a result, the three policies can be jointly optimized while maintaining clear functional specializatio.
3.2 Behavior Cloning
In interactive environments with specialized action and observation spaces, directly training LLM agents with RL often leads to poor sample efficiency. Moreover, since the three policies $\pi^{h}_{\theta}$, $\pi^{l}_{\theta}$ and $\pi^{p}_{\theta}$ must rapidly internalize their respective roles and output structures, we initialize the agent using expert demonstrations via behavior cloning.
We construct three expert demenstration datasets, $\mathcal{D}^{h}$, $\mathcal{D}^{l}$ and $\mathcal{D}^{p}$ based on the Eqs. (3), (4) and (5). Specifically, We index tasks by $i\in[N]$, subtasks by $k\in[K_{i}]$, and within-subtask steps by $t\in[T_{i,k}]$. The datasets are organized as input-target pairs:
$\mathcal{D}^{p}$ $=\Big\{\big((g_{k},a^{k}_{t-1},o^{k}_{t},p^{k}_{t-1}),\;p^{k}_{t}\big)\Big\},$ (6)
$\mathcal{D}^{l}$ $=\Big\{\big((g_{k},p^{k}_{t},o^{k}_{t}),\;a^{k}_{t}\big)\Big\},$ (7)
$\mathcal{D}^{h}$ $=\Big\{\big((u_{i},G_{k},\hat{p}_{k},o^{k+1}_{0}),\;g_{k+1}\big)\Big\}.$ (8)
For notational convenience, we uniformly denote the policy input by $s$ and the action by $u$ across all three policies. Under this unified notation, behavior cloning optimizes each policy by:
$\mathcal{L}_{\mathrm{BC}}(\theta)=-\mathbb{E}_{(s,u)\sim\mathcal{D}}\left[\log\pi_{\theta}(u\mid s)\right],$ (9)
where $\mathcal{D}$ denotes the corresponding expert demonstration dataset for each policy. And the conditional log-likelihood $\log\pi_{\theta}(u\mid s)$ is computed autoregressively. Let $u=(w^{(1)},\ldots,w^{(L)})$ denote the tokenization of the target output and $u^{(<\ell)}=(w^{(1)},\ldots,w^{(\ell-1)})$ denote the preceding tokens. Then:
$\log\pi_{\theta}(u\mid s)=\sum_{\ell=1}^{L}\log\pi_{\theta}\!\left(u^{(\ell)}\mid s,u^{(<\ell)}\right).$ (10)
This behavior cloning procedure serves as an effective initialization for subsequent RL stages. Empirically, even without further RL, our step-level behavior cloning alone achieves superior performance compared to existing baselines, as demonstrated in Section 4.
3.3 Step-Level Offline RL
To further improve generalization, we collect an additional dataset $\tilde{\mathcal{D}}$ based on the behavior-cloned policies for offline optimization. We then combine the collected data with expert demonstrations to form the offline dataset $\mathcal{D}_{\mathrm{r}}=\mathcal{D}\cup\tilde{\mathcal{D}}$, and optimize the policies on $\mathcal{D}_{r}$ using an actor-critic framework. We emphasize that the state $s$ corresponds to a single-step state defined in Eqs. (6)–(8), instead of the full interaction history, which aligns with our step-level formulation.
Utterance-Level Implicit Value Learning.
We implement the critic as an utterance-level value estimator based on the hidden state of the last token. Concretely, given the final-token hidden state $H\in\mathbb{R}^{B\times d},$ the critic attaches two lightweight MLP heads that output scalar predictions for the state-value function $V_{\psi}(s)$ and the action-value function $Q_{\phi}(s,u)$, respectively.
Following the implicit value learning paradigm introduced in Implicit Q-Learning (IQL) and its language adaptation ILQL Kostrikov et al. (2021); Snell et al. (2023), we jointly learn the $Q_{\phi}$ and the $V_{\psi}$ using step-level transitions $(s,u,r,s^{\prime})$ rather than full trajectories. The Q-function is trained by minimizing a TD regression loss bootstrapped from the value function:
$\mathcal{L}_{Q}(\phi)=\mathbb{E}_{(s,u,r,s^{\prime})\sim\mathcal{D}_{\mathrm{r}}}\Big[(r$ $+\gamma V_{\bar{\psi}}(s^{\prime})$ (11)
$-Q_{\phi}(s,u))^{2}\Big],$
where $V_{\bar{\psi}}$ denotes a softly updated target value network used to stabilize training Haarnoja et al. (2018). To approximate the constrained Bellman optimality operator without explicitly maximizing over actions, the value function $V_{\psi}(s)$ is trained using expectile regression. Specifically, $V_{\psi}(s)$ is optimized to regress toward the action-value estimates under an asymmetric squared loss:
$\mathcal{L}_{V}(\psi)=\mathbb{E}_{(s,u)\sim\mathcal{D}_{\mathrm{r}}}\left[L_{2}^{\tau}\big(Q_{\bar{\phi}}(s,u)-V_{\psi}(s)\big)\right].$ (12)
where we define $d=Q_{\bar{\phi}}(s,u)-V_{\psi}(s)$ and $L_{2}^{\tau}(d)=\left|\tau-\mathbf{1}(d<0)\right|d^{2}$ is the expectile loss with expectile parameter $\tau\in(0,1)$ By choosing $\tau>0.5$, the value function approximates an upper expectile of the maximum. This mechanism implicitly biases learning toward high-value actions, enabling stable offline RL without an explicit policy optimization step.
Implicit Policy Improvement via Advantage-Weighted Regression.
The policy is trained to assign higher likelihood to actions with higher estimated value under the learned critic. Concretely, given step-level transitions from the offline dataset $\mathcal{D}_{r}$, the policy is optimized by regressing toward actions favored by the critic. We employ an advantage-weighted regression objective:
$\mathcal{L}_{A}(\theta)=-\mathbb{E}\Big[\exp\!\Big(\frac{A(s,u)}{\beta}\Big)\,\log\pi_{\theta}(u\mid s)\Big],$ (13)
$\text{where}\quad A(s,u)=Q_{\phi}(s,u)-V_{\psi}(s).$
The computation of $\log\pi_{\theta}(u\mid s)$ follows the same autoregressive procedure as in Eq. (10). Following prior offline RL methods Peng et al. (2019); Kostrikov et al. (2021); Nair et al. (2020), we employ an exponential advantage-weighted objective. The temperature parameter $\beta$ controls the sharpness of the weighting and balances policy improvement strength and training stability in our setting.
We apply the offline RL procedure to all three policies, $\pi^{h}_{\theta}$, $\pi^{l}_{\theta}$, and $\pi^{p}_{\theta}$, each equipped with a separate critic network while sharing the same policy parameters. This design provides task-specific value supervision at different levels of abstraction, while shared policy parameters facilitate effective knowledge transfer across hierarchical decisions. Consequently, the model captures complementary decision patterns across levels and achieves more consistent and sample-efficient learning across tasks.
4 Experiments
4.1 Experimental Settings
Benchmarks and Datasets.
We evaluate our approach on two challenging benchmarks:
  • ScienceWorld (Wang et al., 2022a) is a text-based interactive benchmark with 30 science task families (e.g., physics, chemistry, biology), each containing many parameterized variants, yielding hundreds to thousands of tasks that require multi-step reasoning and experimentation.
  • ALFWorld (Shridhar et al., 2020) is a household task benchmark aligned with ALFRED, comprising134 language-conditioned tasks across 6 task types (e.g., pick-and-place, cleaning, heating), focusing on long-horizon action planning.
For dataset construction, we generate hierarchical annotations, including subtasks and local progress signals from expert trajectories using DeepSeek. The prompts used for subtask decomposition and progress generation are provided in the Appendix A. For offline RL, we collect additional trajectories using policies initialized via behavior cloning. Following the experimental setup of GLIDER Hu et al. (2025), we adopt a trajectory mixture ratio of 1:21\!:\!2, which was identified as the most effective setting in their study.
Models and Baselines.
We evaluate STEP-HRL on three outstanding open source models: Mistral-7B Jiang et al. (2023), Gemma-7B Team et al. (2024) and Llama3-8B Meta AI (2024).
We compare against the following baselines: 1) ReAct (Yao et al., 2022), a prompting framework that interleaves reasoning traces and environment actions in a Thought–Action–Observation loop. 2) Reflexion (Shinn et al., 2023), which improves subsequent trials by storing self-reflective feedback in an episodic memory. 3) SwiftSage (Lin et al., 2023), a dual-process agent that combines a behavior-cloned action model with an LLM-based planner for interactive tasks. 4) ETO (Song et al., 2024), which iteratively collects contrastive (failure/success) trajectories and optimizes the policy via DPO Rafailov et al. (2023). 5) WKM (Qiao et al., 2024), which augments planning with a parametric world knowledge model that provides task priors and dynamic state knowledge. 6) GLIDER (Hu et al., 2025), an offline HRL framework that decomposes complex tasks and learns high-level and low-level policies for decision making. We also report results of ChatGPT (gpt-3.5-turbo-0125) and GPT-4 (gpt-4-32k-0613) for comparison by referencing previously published results (Qiao et al., 2024).
Training Details.
All fine-tuning baselines and our method are fine-tuned using LoRA (Hu et al., 2022). For behavior cloning, we train the policies for 5 epochs with a learning rate of $1\times 10^{-4}$ and a batch size of 128. During the offline RL stage, we train for 3 epochs, using learning rates of $1\times 10^{-5}$ and $1\times 10^{-4}$ for the actor and critic, respectively, with a batch size of 256. All models are optimized using AdamW (Loshchilov and Hutter, 2017) optimizer. All experiments are conducted on 8 NVIDIA A100 80G GPUs. Detailed hyperparameters and additional experimental settings are provided in Appendix  B.
4.2 Results
Table 1: Main Results. We report performance comparisons for three backbone models on the ScienceWorld and ALFWorld benchmarks. corresponds to prompt-based methods without updating model parameters, and corresponds to LoRA-based fine-tuning approaches. $\uparrow$ indicates the relative improvement of STEP-HRL compared to the best results among the baselines.
Backbone Method ScienceWorld ALFWorld
Seen Unseen Seen Unseen
GPT-3.5-Turbo ReAct 8.57 5.97 15.41 13.99
GPT-4 44.29 38.05 67.32 65.09
Mistral-7B ReAct 20.72 17.65 7.86 5.22
Reflexion 21.07 18.11 11.56 6.00
SwitchSage 48.40 45.25 30.29 26.52
ETO 58.17 51.85 66.84 71.43
WKM 62.12 53.62 73.57 76.87
GLIDER 67.31 65.14 70.02 74.83
STEP-HRL 80.28 ($\uparrow 19.27\%$) 75.21 ($\uparrow 15.46\%$) 96.43 ($\uparrow 31.07\%$) 97.01 ($\uparrow 26.20\%$)
Gemma-7B ReAct 3.58 3.51 6.43 2.24
Reflexion 4.94 3.93 7.14 2.99
SwitchSage 33.43 30.90 8.23 5.72
ETO 50.44 47.84 66.43 68.66
WKM 53.68 49.24 70.71 70.40
GLIDER 63.67 58.50 72.12 70.88
STEP-HRL 78.89 ($\uparrow 24.02\%$) 74.08 ($\uparrow 26.63\%$) 97.86 ($\uparrow 35.69\%$) 97.76 ($\uparrow 37.92\%$)
Llama-3-8B ReAct 24.76 22.66 2.86 3.73
Reflexion 27.23 25.41 4.29 4.48
SwitchSage 42.22 40.58 20.39 10.78
ETO 57.90 52.33 64.29 64.18
WKM 60.12 54.75 68.57 65.93
GLIDER 77.43 68.34 71.56 75.38
STEP-HRL 81.57 ($\uparrow 5.35\%$) 77.81 ($\uparrow 13.86\%$) 97.14 ($\uparrow 35.75\%$) 97.76 ($\uparrow 29.69\%$)
Main results.
Table 1 reports the evaluation results of STEP-HRL across three backbone models on the ScienceWorld and ALFWorld benchmarks. Across all settings, STEP-HRL consistently outperforms strong prior baselines on both seen and unseen tasks. On ALFWorld, STEP-HRL achieves near-saturated performance, with success rates exceeding 90% across different backbone models. On ScienceWorld, STEP-HRL also yields consistent and substantial improvements over existing methods, demonstrating its effectiveness in more challenging and diverse environments. Notably, the performance gap between backbones is substantially reduced, indicating strong robustness and scalability of our proposed framework.
Performance across Different Model Scales.
Table 2: Performance of STEP-HRL on ScienceWorld and ALFWorld across model scales.
Model ScienceWorld ALFWorld
Seen Unseen Seen Unseen
Llama-1B 51.88 49.78 89.86 89.60
Llama-3B 65.31 61.79 94.29 94.00
Llama-8B 80.57 77.81 96.43 97.76
Table 2 summarizes the performance of STEP-HRL across different model scales. Overall, performance improves steadily as model capacity increases, with larger backbones achieving better performance. Notably, even smaller models such as Llama-1B and Llama-3B demonstrate competitive performance, particularly on ALFWorld. This trend suggests that STEP-HRL is exceptionally effective even across a wide range of model scales, while additional model capacity further enhances robustness and generalization, especially on more challenging ScienceWorld tasks.
4.3 Ablation Studies
As shown in Figure 2, we examine the effectiveness of key components in STEP-HRL. We consider three variants: 1) w/o LP, which removes the local progress policy, forcing the low-level and high-level policies to condition on $(g_{k},o^{k}_{t})$ and $(c_{i},G_{k},o^{k+1}_{0})$ respectively; 2) w/o Hier, which eliminates the hierarchical structure and directly relies on the local progress module to summarize the global interaction history; 3) w/o RL, which omits the offline RL stage and reduces training to behavior cloning only. All variants are trained using the same step-level data for a fair comparison.
Across all settings, alternative variants consistently lead to degraded performance. Most notably, the local progress module plays a central role by condensing subtask-relevant interaction information into a compact summary Without this module, many states become indistinguishable, making credit assignment and policy optimization significantly more challenging. The hierarchical structure further contributes by decomposing complex tasks into manageable subtasks, which alleviates the burden on the local progress module and prevents it from being overwhelmed when summarizing long-horizon interactions. Finally, the offline RL stage refines the policies beyond behavior cloning, improving generalization to unseen tasks through value-guided policy updates. Collectively, these results highlight the complementary roles of all components and underscore the importance of the proposed design in STEP-HRL.
Figure 2: Ablation study of STEP-HRL on unseen ScienceWorld tasks with different backbone models. w/o LP denotes removing the local progress policy, w/o Hier denotes removing the hierarchical structure, and w/o RL denotes removing the offline RL stage, reducing the training procedure to behavior cloning only.
4.4 Analysis on Efficiency
Token Usage Efficiency.
As shown in Figure 3, we analyze the token efficiency of STEP-HRL in comparison with standard RL (i.e., a normal ReAct-style agent) and HRL on an ALFWorld task. For a fair comparison, all methods are evaluated under the same observation and action sequence. The task is decomposed into four subtasks and requires a total of 29 environment steps to complete.
Standard RL incurs steadily increasing per-step token costs as the interaction progresses, since the policy repeatedly conditions on an ever-growing interaction history. Although HRL reduces the token usage by decomposing the task into subtasks, it exhibits substantial variability, with pronounced spikes at subtask generation steps where long accumulated contexts are processed. Such high variance in input sequence not only increases inference latency, but also leads to inefficient training. In particular, GPU memory allocation must accommodate peak input lengths induced by high-level samples, resulting in underutilization during most steps and reduced overall training efficiency.
In contrast, STEP-HRL maintains an approximately constant per-step token usage. By leveraging compact summaries of both global task progress and local subtask progress, STEP-HRL avoids conditioning on full interaction histories. This design effectively bounds the per-step inference cost, yielding the lowest average token usage with minimal variance. Overall, the result highlights the advantage of STEP-HRL in enabling predictable, efficient and well-balanced inference and training behavior, which is particularly important for long-horizon interactive environments.
Wall-Clock Inference Efficiency.
We further conduct analysis on wall-clock inference latency. As shown in Table 3, STEP-HRL is slightly slower than normal agent framework in early steps. This is because STEP-HRL contains three policies and therefore performs more forward passes per step. Under KV-cache decoding, this introduces additional prefills, and local-progress generation typically produces more tokens than primitive actions.
However, as the interaction history grows, the latency of the history-conditioned baseline increases, whereas STEP-HRL remains stable. This suggests that STEP-HRL is less sensitive to history length and scales more favorably in long-horizon settings.
Table 3: Per-step inference latency on a single A100 GPU in the ALFWorld Pick&Place. $\star$ denotes subtask generation, which introduces an additional forward pass.
Step Ours (s) Normal (s)
1 1.8 $\pm$ 0.3 1.3 $\pm$ 0.3
10 1.6 $\pm$ 0.2 1.4 $\pm$ 0.2
20 1.5 $\pm$ 0.2 1.6 $\pm$ 0.2
30 1.5 $\pm$ 0.3 1.7 $\pm$ 0.3
40 1.6 $\pm$ 0.2 1.9 $\pm$ 0.2
Figure 3: Simulated per-step token usage on ALFWorld pick_two_obj_and_place task under identical observation and action sequence across three RL paradigms.
4.5 Analysis on Offline RL
Figure 4 presents a sensitivity analysis of the proposed offline RL procedure with respect to key factors. We observe that the advantage temperature $\beta$ plays a critical role in balancing update aggressiveness and stability: among the tested values, $\beta=0.95$ consistently yields the highest final performance, while both smaller and larger temperatures lead to inferior convergence. Similarly, varying the expectile parameter $\tau$ reveals that a higher expectile (e.g., $\tau=0.9$) provides more effective value estimation and results in stronger policy improvement compared to lower settings.
We further analyze the impact of data sources used for offline RL. Training on mixed datasets that combine expert demonstrations with BC-collected trajectories consistently outperforms using either expert-only or BC-collected data alone. While BC-collected data contains informative failure trajectories that are valuable for policy improvement, it also introduces lower-quality and noisier samples compared to expert demonstrations. We also study the effect of scaling the amount of training data. Increasing data size improves performance up to a point, with twice the expert data achieving the best overall results. Beyond this regime, additional data yields diminishing returns, as excessively large datasets introduce redundant or low-quality samples that hinder stable learning.
Figure 4: Offline RL sensitivity analysis with respect to advantage temperature $\beta$, expectile parameter $\tau$, data sources, and data amounts across training.
5 Related Work
LLM Agents.
With their strong semantic understanding and emergent reasoning abilities, large language models (LLMs) have been explored as autonomous agents for decision making in complex and interactive environments Guo et al. (2024); Wang et al. (2024). Early studies primarily adopt prompt-based formulations, where agents generate intermediate reasoning traces to support multi-step decisions, such such as Chain-of-Thought Wei et al. (2022), ReAct Yao et al. (2022), Reflexion Shinn et al. (2023) and their variants Yao et al. (2023); Wang et al. (2022b). Subsequent work augments LLM agents with additional system components, including tool use Schick et al. (2023); Qin et al. (2023); Wu et al. (2024); Li et al. (2025), memory mechanisms Zhang et al. (2025); Sarch et al. (2024); Xu et al. (2025), and multi-agent coordination Chen et al. (2024a); Bo et al. (2024); Estornell et al. (2024). Beyond architectural augmentation, another line of work focuses on grounding LLM agents through learning from expert demonstration via fine tuning (behavior cloning), demonstrating strong gains  Zeng et al. (2024b); Chen et al. (2023); Yin et al. (2023); Chen et al. (2024b). However, these approaches heavily rely on high-quality expert data and trajectory-level supervision, and their performance degrades in long-horizon decision making and complex interactive tasks due to limited exploration and severe distribution shift..
Reinforcement Learning in LLM Agents.
Reinforcement learning (RL) has achieved notable success in aligning and improving large language models Ouyang et al. (2022); Shao et al. (2024), and has also proven effective for training LLM agents through explicit reward and penalty mechanisms. Most prior work adopts an interaction-driven pipeline, where an LLM agent receives goal-directed feedback from the environment and is fine-tuned using RL algorithms such as PPO Schulman et al. (2017); Zhai et al. (2024); Szot et al. (2023); Peiyuan et al. (2024). Preference-based methods, such as ETO Song et al. (2024), further collect contrastive trajectory pairs from environment interactions and update LLM policies via preference optimization objectives like DPO Rafailov et al. (2023). To achieve fine-grained reinforcement learning signals instead of optimizing the full trajectories, Wen et al. (2024) decompose RL objectives to provide action-level feedback for LLM agents, while GiGPO Feng et al. (2025) hierarchically estimates step-level advantages to improve training efficiency. Despite their effectiveness, most existing RL-based LLM agents depend on full interaction histories for decision making, where the increasing context length poses significant challenges for credit assignment and computational efficiency.
To mitigate this limitation, several works have explored HRL frameworks such as EPO Zhao et al. (2024) and GLIDER Hu et al. (2025) decompose complex tasks into subtasks and learn coordinated high-level and low-level policies. However, even docompsing or fine-grained optimization, these methods still rely on history-conditioned policies, resulting in inefficient credit assignment and high computational overhead.
6 Conclusion
In this paper, we proposed STEP-HRL, a innovative framework that enables efficient step-level learning for LLM agents without relying on full interaction histories. STEP-HRL decomposes tasks into a hierarchical structure and introduces a local progress module to summarize subtask-relevant information, allowing both high-level and low-level policies to operate on compact, step-level state representations. Empirical results on ScienceWorld and ALFWorld demonstrate that STEP-HRL consistently improves performance and generalization. Overall, STEP-HRL provides a practical and scalable approach for training LLM agents. We believe that step-level abstraction with structured progress summaries offers a promising direction for improving both efficiency and robustness in future LLM agent research.
Limitations
Despite the effectiveness of STEP-HRL, it still has several limitations worth noting:
  • STEP-HRL highly relies on high-quality expert demonstrations. In particular, the construction of step-level data requires carefully designed subtask and local progress. Designing and curating them can be non-trivial in practice, especially for complex environments with ambiguous subtask boundaries or poorly defined progress signals. This reliance may limit the applicability of STEP-HRL in domains where expert data or structured supervision is scarce.
  • In our implementation, subtask termination is predicted jointly with primitive actions, such that each low-level output includes both an action and a termination indicator. This design may result in inaccurate termination decisions, including premature termination or delayed subtask completion. Such errors can degrade the quality of collected transitions and introduce bias into critic value estimation, which in turn may cause misalignment between high-level planning and low-level execution during inference.
Acknowledgments
The work was supported by the National Natural Science Foundation of China (Grant No. U22B2019), China Scholarship Council (No. 202406470057) and Beijing University of Posts and Telecommunications-China Mobile Communications Group Co., Ltd. Joint Institute.
Appendix A Benchmarks and Datasets
A.1 Benchmarks
We evaluate our approach on two widely used language-based interactive decision-making benchmarks: ScienceWorld and ALFWorld.
  • ScienceWorld (Wang et al., 2022a) is a text-based environment designed for science experimentation. It consists of 30 tasks spanning 10 categories, where agents are required to demonstrate scientific reasoning through interactive exploration. The environment provides dense rewards at each step, with values ranging from 0 to 1, reflecting incremental task progress.
  • ALFWorld (Shridhar et al., 2020) simulates household environments that involve navigation and object manipulation. In contrast to ScienceWorld, ALFWorld adopts a sparse reward setting, where an agent receives a reward of 1 only upon successful task completion and 0 otherwise.
Both ScienceWorld and ALFWorld are evaluated under two settings: Seen and Unseen. The Seen split contains in-distribution tasks that follow the similar task and variations as those observed during training, and is used to evaluate in-distribution performance. In contrast, the Unseen split consists of out-of-distribution task variations with novel mechanism or object, and is used to assess the generalization ability. Dataset statistics for all splits are summarized in Table 4.
Table 4: Dataset statistics.
Dataset Train Seen Unseen
ScienceWorld 1,483 194 211
ALFWorld 3,211 140 134
A.2 Datasets
Expert Dataset.
For subtask and local-progress generation, we employ a combination of rule-based heuristics and the DeepSeek model. For ScienceWorld, due to the substantial structural diversity across tasks, we adopt task-specific prompts and subtask decomposition strategies tailored to each task category. In contrast, for ALFWorld, we design a unified prompt that guides DeepSeek to generate both subtask and local-progress fields in a consistent manner. Figure 8 presents the prompt used for generating local progress.
After generating subtask and local progress fields, we construct the SFT (BC) datasets for all three policies. The data structure of these datasets are shown below, and the prompts used during training and inference are provided in Figure 6.
Training Dataset Structure High-Level SFT Data:
Input
: { high prompt, task description, current observation, completed subtasks, previous local progress }
Target: next subtask Low-Level SFT Data:
Input
: {low prompt, subtask, current observation, local progress}
Target: action Local-Progress SFT Data:
Input
: {local progress prompt, subtask, executed action, resulting observation, previous local progress}
Target: updated local progress
Offline RL Dataset.
We construct the RL dataset by combining expert demonstrations with trajectories collected from behavior-cloned policies.
During data collection, we adopt different sampling temperatures for different policy components to balance exploration and action validity. Specifically, for the high-level policy and the local-progress policy, we set the sampling temperature to 1.0 to encourage diverse subtask sequences and reasoning paths. In contrast, the low-level policy generates primitive actions that must strictly conform to the environment’s action format and input constraints. To avoid producing invalid or malformed actions, we therefore set the sampling temperature of the low-level policy to 0, ensuring deterministic and well-formed action generation.
The final offline RL dataset consists of both expert data and policy-collected trajectories, with an approximate ratio of 1:2. Among the collected trajectories, around 25% correspond to unsuccessful episodes. Including such unsuccessful data enables the model to observe negative outcomes and learn to penalize suboptimal actions, which is beneficial for stable offline policy learning.
Local Progress Prompt (DeepSeek-ALFWorld) You are an AI agent responsible for updating local progress, a short cumulative summary of what has been achieved within the current subtask. Inputs: current subtask previous local progress current action observation Global Rules: 1. Exact token matching: All object and location names MUST EXACTLY match strings in the subtask or observation. Do NOT rephrase or normalize names. 2. No invented facts: Do NOT infer properties, conditions, or failures unless explicitly stated. 3. Task type: Subtasks starting with Locate and pick up or Locate and use are Locate tasks; all others are Non-Locate tasks. Locate Tasks Output: <progress sentence> $|$ [Checked: loc1, loc2, …] 1. Checked update: Retain all previous locations. Add the current location ONLY IF a search action is taken and the target is confirmed NOT present. 2. Termination: If the target is found or picked up, the search ends and must not continue. 3. New + Except: For new <OBJ> except <OBJ> N, any different-numbered <OBJ> completes the subtask immediately. 4. Unsuccessful search: The sentence must include unchecked and imply door states (opened / closed), without mentioning specific unchecked locations. 5. Language constraints: Do NOT mention checked locations, reuse fixed templates, or repeat more than 3 consecutive words from the previous progress. Non-Locate Tasks (Place / Clean / Heat / Cool / Use) Output: <progress sentence> 1. Do NOT include Checked or any tags. Mention the object ONLY IF it appears in the current action. 2. No existence or location statements: Do NOT state or imply object presence, absence, containment, or discovery. 3. No failure reasoning: Do NOT explain progress via unmet conditions or missing objects. 4. Assumed availability: Treat the target object as available by definition of the subtask. 5. Allowed content only: Describe only the executed operation or its direct state change.
Figure 5: Full prompt specification for local progress annotation on the ALFWorld expert dataset (DeepSeek).
Prompt in Training and Inference High-Level Prompt:
You are a high-level planner. Based on the state (task description, historical subtasks, last subtask progress and current observation), please generate a clear and simple subtask.
Low-Level Prompt:
You are a low-level action executor. Based on the current subtask, observation and local progress, please generate a executable action and determine if the subtask is completed (True/False).
Local-Progress Prompt:
You are an AI agent responsible for updating local progress within a subtask. Based on the current subtask, the previous local progress, the current action, and the resulting observation, update the local progress.
Figure 6: The prompt used in training and inference stages.
Appendix B Training Details
Models.
We conduct our main experiments using the following instruction-tuned large language models:
  • mistralai/Mistral-7B-Instruct-v0.2
  • google/gemma-1.1-7b-it
  • meta-llama/Meta-Llama-3-8B-Instruct
For scalability experiments across different model sizes, we additionally evaluate:
  • meta-llama/Llama-3.2-1B-Instruct
  • meta-llama/Llama-3.2-3B-Instruct
Hyperparameters.
The details of all hyperparameters are summarized in Table 5. We adopt LoRA for parameter-efficient fine-tuning across all models. During the behavior cloning (BC) stage, models are trained for 5 epochs using the AdamW optimizer with a learning rate of $1\times 10^{-4}$. Training is performed with a total batch size of 128, achieved via a per-device batch size of 8 and 2 gradient accumulation steps, which balances computational efficiency and training stability.
For the offline reinforcement learning stage, we train the model for 3 epochs, using separate learning rates for the actor ($1\times 10^{-5}$) and critic ($1\times 10^{-4}$) . The target critic is updated via soft updates with coefficient $\tau_{1}=0.2$, and the discount factor is set to $\gamma=0.99$. To further stabilize training, we first warm up the critic for 100 steps before jointly optimizing the actor and critic. We employ an advantage-weighted objective with weighting factor $\beta=0.95$, and adopt expectile regression with parameter $\tau_{2}=0.95$ for value learning.
During trajectory sampling, we use a sampling temperature of 0.7 for the high-level and local-progress policies to encourage diverse reasoning paths, while setting the temperature to 0 for the low-level policy to ensure deterministic and valid primitive action generation. We further constrain text generation by allowing a maximum of 32 tokens for the high-level and low-level policies, and 150 tokens for the local-progress policy.
For evaluation, we impose a maximum of 50 environment steps per episode for both ALFWorld and ScienceWorld tasks, ensuring a consistent evaluation budget across benchmarks.
Table 5: STEP-HRL hyperparameters.
Hyperparameter Value
Optimization
batch size 128
batch size per device 8
gradient accumulation steps 2
optimizer AdamW
actor learning rate $1\times 10^{-5}$
critic learning rate $1\times 10^{-4}$
sft learning rate $1\times 10^{-4}$
Training schedule
sft epochs 5
orl epochs 3
orl warmup steps 100
RL-specific
discount factor $\gamma$ 0.99
advantage weighted factor $\beta$ 1
soft update $\tau_{1}$ 0.2
expectile parameter $\tau_{2}$ 0.95
Generation
sampling temperature 0.7
max new tokens ($\pi_{\theta}^{h}$, $\pi_{\theta}^{l}$) 32
max new tokens ($\pi_{\theta}^{p}$) 150
LoRA
lora $r$ 16
lora alpha 32
lora dropout 0.05
lora target modules q_proj, k_proj,
v_proj, o_proj
Data
data mixture ratio 1:2
env limit steps 50
Appendix C Evaluation Details
We further report model performance on each individual task family. Since the result distributions are similar across different backbone models, we only present results for Llama-3-8B in Tables 6 and 7.
Table 6: Evaluation details on ALFWorld unseen task.
Task ID Task Name #Variants Success Rate (%) Avg. Steps (Succ.) Avg. Steps (All)
1 Pick&Place 24 100.0 12.7 12.7
2 Examine in Light 18 100.0 13.3 13.3
3 Clean&Place 31 100.0 10.6 10.6
4 Heat&Place 23 100.0 17.1 17.1
5 Cool&Place 21 100.0 15.7 15.7
6 Pick Two&Place 17 82.4 21.8 26.6
Total 134 97.8 14.7 15.3
Table 7: Evaluation details on ScienceWorld unseen task.
Task ID Task Name #Variants Avg Score Avg. Steps (Succ.) Avg. Steps (All)
0 boil 9 68.9 45.0 48.3
1 change-the-state-of-matter-of 9 62.2 34.0 46.4
2 chemistry-mix 8 67.8 20.8 25.5
3 chemistry-mix-paint-secondary-color 9 88.9 9.5 9.4
4 chemistry-mix-paint-tertiary-color 9 54.4 16.8 13.7
5 find-animal 10 100.0 11.6 11.6
6 find-living-thing 10 100.0 11.6 11.6
7 find-non-living-thing 10 100.0 5.8 5.8
8 find-plant 10 97.5 10.0 14.0
9 freeze 9 55.0 29.0 42.6
10 grow-fruit 10 43.4 46.4
11 grow-plant 10 98.8 34.8 35.4
12 identify-life-stages-1 5 77.0 25.0 27.8
13 identify-life-stages-2 4 25.0 5.0 6.2
17 lifespan-longest-lived 10 100.0 4.0 4.0
18 lifespan-longest-lived-then-shortest-lived 10 100.0 5.0 5.0
19 lifespan-shortest-lived 10 100.0 4.0 4.0
20 measure-melting-point-known-substance 10 39.1 28.5
22 melt 9 61.9 34.7 39.8
25 power-component 5 100.0 11.8 11.8
26 power-component-renewable-vs-nonrenewable-energy 5 21.2 37.8
27 test-conductivity 10 78.3 16.4 19.7
28 test-conductivity-of-unknown-substances 10 70.0 17.1 18.7
29 use-thermometer 10 100.0 17.6 17.6
Total 211 77.8 15.45 22.0
Appendix D Case Study
Trajectory Comparison Normal Task Description: Your task is to find a(n) non-living thing. First, focus on the thing. Then, move it to the purple box in the bathroom. Observation: This room is called the kitchen. In it, you see the agent, a substance called air, a chair, a counter, a cupboard, a freezer, a fridge, a glass jar, a lighter, an oven, a painting, a sink, soap, a stopwatch, a stove, a table with a glass cup, and a thermometer. You also see: a door to the bathroom (closed), a door to the hallway (closed), and a door to the outside (closed). Action: open door to bathroom Reward: $0.08$ Observation: The door is now open. Action: go to bathroom Reward: $0.17$ Observation: This room is called the bathroom. In it, you see a bathtub, a drawing, a glass cup, a purple box, a sink, and a toilet. Action: focus on cup Reward: $0.50$ Observation: You focus on the glass cup. Action: move cup to purple box Reward: $0.25$ Observation: You move the glass cup to the purple box. done. STEP-HRL Task Description: Your task is to find a(n) non-living thing. First, focus on the thing. Then, move it to the purple box in the bathroom. Observation: This room is called the kitchen. In it, you see the agent, a substance called air, a chair, a counter, a cupboard, a freezer, a fridge, a glass jar, a lighter, an oven, a painting, a sink, soap, a stopwatch, a stove, a table with a glass cup, and a thermometer. You also see: a door to the bathroom (closed), a door to the hallway (closed), and a door to the outside (closed). Subtask 1: Navigate to the bathroom and focus on any non-living object. Action: open door to bathroom Reward: $0.08$ Observation: The door is now open. Local Progress: Agent is in the kitchen and the door to the bathroom is open, not yet in the bathroom. [Route: kitchen] Action: go to bathroom Reward: $0.17$ Observation: This room is called the bathroom. In it, you see a bathtub, a drawing, a glass cup, a purple box, a sink, and a toilet. Local Progress: Agent is now in the bathroom, ready to search for a non-living thing. [Route: kitchen $\rightarrow$ bathroom] Action: focus on cup Reward: $0.50$ Observation: You focus on the glass cup. Local Progress: Agent has found and focused on a glass cup, a non-living thing, in the bathroom. [Route: kitchen $\rightarrow$ bathroom] Subtask 2: Move the non-living thing to the purple box in the bathroom. Action: move cup to purple box Reward: $0.25$ Observation: You move the glass cup to the purple box. done.
Figure 7: Case study comparing a standard ReAct-style agent and STEP-HRL on a ScienceWorld task. Normal paradigms make decisions based on the accumulated interaction trajectory, which tends to collapse in long-horizon complex tasks. In contrast, STEP-HRL makes decisions based solely on the current observation, augmented with high-level subtask and step-wise local progress, enabling step-level decision making.
STEP-HRL in Long-Horizon Complex Task Task Description: Your task is to boil water. For compounds without a boiling point, combusting the substance is also acceptable. First, focus on the substance. Then, take actions that will cause it to change its state of matter. Observation: You are in art studio. You also see: A door to the hallway (that is closed). Subtask 1: Navigate to kitchen
Action:
open door to hallway   Reward: 0.000.00Observation: The door is now open.   Local Progress: Navigating to kitchen and the door to hallway is now open.   Action: go to hallway   Reward: 0.000.00Observation: You move to the hallway.   Local Progress: Moving from art studio to hallway [Route: art studio -> hallway]Action: open door to kitchen   Reward: 0.000.00Observation: The door is now open.   Local Progress: The kitchen door is open and ready to go. [Route: art studio -> hallway]Action: go to kitchen   Reward: 0.000.00Observation: You move to the kitchen.   Local Progress: Arrived in the kitchen. [Route: art studio -> hallway -> kitchen]
Subtask 2: Prepare tools for measuring temperature and boiling
Action:
pick up thermometer   Reward: 0.000.00Observation: You move the thermometer to the inventory.   Local Progress: Thermometer obtained.   Action: open cupboard   Reward: 0.000.00Observation: The cupboard is now open.   Local Progress: Thermometer obtained; cupboard opened, revealing a metal pot and other items.   Action: pick up metal pot   Reward: 0.000.00Observation: You move the metal pot to the inventory.   Local Progress: Metal pot obtained; thermometer already in inventory.
Subtask 3: Fill metal pot with water
Action:
move metal pot to sink   Reward: 0.000.00Observation: You move the metal pot to the sink.   Local Progress: The metal pot is now in the sink, ready to be filled.   Action: activate sink   Reward: 0.030.03Observation: The sink is now activated.   Local Progress: The metal pot is now filled with water in the sink.   Action: deactivate sink   Reward: 0.000.00Observation: The sink is now deactivated.   Local Progress: The metal pot remains filled with water in the sink, which is now turned off.   Action: pick up metal pot   Reward: 0.000.00Observation: You move the metal pot to the inventory.   Local Progress: The metal pot containing water is now in the agent’s inventory.
Subtask 4: Focus on substance
Action:
focus on substance in metal pot   Reward: 0.670.67Observation: You focus on the water.   Local Progress: You focus on the substance called water in the metal pot.
Subtask 5: Heat water with stove in kitchen
Action:
move metal pot to stove   Reward: 0.020.02Observation: You move the metal pot to the stove.   Local Progress: Metal pot containing water is now on the stove.   Action: activate stove   Reward: 0.000.00Observation: The stove appears broken, and can’t be activated or deactivated.   Local Progress: The stove is broken and cannot be activated; the metal pot with water remains on it.
Subtask 6: Navigate to foundry to heat water because stove is broken
Action:
pick up metal pot   Reward: 0.000.00Observation: You move the metal pot to the inventory.   Local Progress: Agent has picked up the metal pot containing water.   Action: open door to outside   Reward: 0.000.00Observation: The door is now open.   Local Progress: Agent is still in the kitchen with the pot of water, and the door to outside is open. [Route: kitchen]Action: go to outside   Reward: 0.000.00Observation: You move to the outside.   Local Progress: Agent is now outside, heading toward the foundry. [Route: kitchen -> outside]Action: open door
to foundry   Reward: 0.000.00Observation: The door is now open.   Local Progress: Agent has opened the foundry door but has not yet entered. [Route: kitchen -> outside]Action: go to foundry   Reward: 0.000.00Observation: You move to the foundry.   Local Progress: Agent has entered the foundry. [Route: kitchen -> outside -> foundry]
Subtask 7: Heat water with blast furnace
Action:
open blast furnace   Reward: 0.000.00Observation: The blast furnace is now open.   Local Progress: The blast furnace is open and ready to be used.   Action: move metal pot to blast furnace   Reward: 0.010.01Observation: You move the metal pot to the blast furnace.   Local Progress: The metal pot containing water is now inside the open blast furnace.   Action: activate blast furnace   Reward: 0.020.02Observation: The blast furnace is now activated.   Local Progress: The blast furnace is activated and heating the metal pot containing water.
Subtask 8: Monitor water temperature
Action:
examine substance in metal pot   Reward: 0.020.02Observation: a substance called water   Local Progress: Examined water in the pot; temperature not yet measured.   Action: use thermometer in inventory on substance in metal pot   Reward: 0.000.00Observation: the thermometer measures a temperature of 50 degrees celsius   Local Progress: Water temperature measured at 50°C.   Action: examine substance in metal pot   Reward: 0.000.00Observation: a substance called water   Local Progress: Water temperature measured at 50°C.   Action: wait1   Reward: 0.230.23Observation: You decide to wait for 1 iterations.
done.
Figure 8: The trajectory of STEP-HRL on ScienceWorld boil task.